OpenStudy (anonymous):

If the legs of a right triangle are 8 and 10, what is the length of the hypotenuse?

OpenStudy (anonymous):

pythagorean theorem, a^2 + b^2 = c^2

OpenStudy (anonymous):

Use the Pythagorean Theorem a^2 + b^2 = c^2 8^2 + 10^2 = c^2

OpenStudy (anonymous):

64 + 100 = c^2

OpenStudy (anonymous):

thanks, srry x.x was busy